Wesley thinks his heartrate will increase as he increases his running speed. To see if this relationship exists, he records eight different speeds and models it with a scatterplot and regression output.

A scatterplot is shown with the x axis labeled Speed, miles per hour, and the y axis labeled Pulse, beats per minute. The points on the graph increase from 0.75 and 72 to 4.1 and 131.
Regression Statistics
Multiple R 95.70928%
R Square 91.60266%
Adjusted R Square 90.20310%
Standard Error 6.513645838
Observations 8
df SS MS F
Regression 1 2,776.934507 2,776.934507 65.45116
Residual 6 254.5654926 42.4275821
Total 7 3031.5
Coefficients Standard Error t Stat p-Value
Intercept 63.06927886 5.643282415 11.17599195 3.06E-05
Speed 18.57636597 2.296159651 8.090189183 0.000191
Coefficients Standard Error t Stat p-Value
Intercept 63.06927886 5.643282415 11.17599195 3.06E-05
Speed 18.57636597 2.296159651 8.090189183 0.000191

Part A: Write the equation of the regression line using the regression output.

Part B: What do the slope and intercept parameters mean using the context of the problem?

Part C: Compute the margin of error that Wesley should use if he wants to provide a 98% confidence interval for the slope. Assume that conditions for inference are satisfied.
Answer #1

A.
Regression equation is
ycap = 63.0693 + 18.5764x

B.
Here slope = 18.5764, indicates the increase in heart rate for every 1 unit increase in the speed of running.
intercept = 63.0693 indicates the base heart-rate, this means even if someone is not running the heart-rate would be 63.0693

C.
For 98%, t- value = 2.998

ME = t*SE
ME = 2.998*18.5764 = 55.6920